### v1

# A simple explanation for the absence of the spherical nuclei

Abstract
Inspired by the EMC effect, the Cd puzzle and the SU3-IBM, a hypothesis can be given for a nucleus, that only the nucleus itself is a trivial (0,0) representation of the SU(3) group, which leads to a simple conclusion that spherical nucleus does not exist and the spherical mean field is not allowed. The key conclusion is that the color degree of freedom is required even at the low energy excitation of the nucleus.
keywords:
Cd puzzle; EMC effect; SU3-IBM; spherical nuclei
In the traditional wisdom, nuclei with magic numbers are spherical [1, 2], because the magic numbers can be explained by the spherical three-dimensional harmonic oscillator potential with strong spin-orbital coupling [3]. This mean-field view, or the shell model, forms the basis of the nuclear structure [2]. When the valence nucleons increase (the number is still small), the spherical surface is still kept, and the residual two-body interaction is expected to play a dominant role, which leads to the phonon vibrational excitation for the even-even nucleus [4, 5].
The Cd nuclei have long been considered as typical example with spherical vibrational excitation [5]. The reason is that the low-lying levels of these nuclei look very similar to the phonon excitation spectra. However the Cd nuclei exist shape coexistence [6, 7], which hinders our understanding of the normal states of the Cd isotopes. It was long recognized that the electromagnetic transitions in the normal states are different from the ones in the spherical phonon mode. Since researchers in the field of nuclear structure believe that the Cd nuclei must be spherical, this deviation is suggested to result from the strong coupling between the normal states and intruder states [8, 9, 10].
No one believes that there are other possibilities, however the Cd puzzle experiments denied this idea, which is a very unexpected result [11, 12, 13, 14, 15, 16]. The coupling between the normal states and the intruder states is in fact weak [13], and the spherical phonon excitation of the Cd nuclei is refuted [4, 5, 17]. The normal states are a new spherical-like $\gamma$-soft collective mode [18, 19, 20]. This means that the origin of the collective nature in the nucleus requires an alternative explanation. Just as Heyde and Wood said: “The emerging picture of nuclear shapes is that quadrupole deformation is fundamental to achieving a unified view of nuclear structure.” and “ a shift in perspective is needed: sphericity is a special case of deformation.” [4]
Inspired by the Cd puzzle, I proposed the extended interacting boson model [21, 22, 23] with SU(3)-higher-order interactions (SU3-IBM) [18], just satisfying the idea of Heyde and Wood. In this model, only the U(5) limit and the SU(3) limit are included. It suggests that the SU(3) symmetry dominates the onset of the quadrupole deformation [24, 25]. Not only the quadrupole deformation but also the SU(3) symmetry is important for the unified description of nuclear structure. The Cd puzzle implies a new spherical-like deformation [18, 19, 20].
In this new spherical-like $\gamma$-soft rotation, it looks very similar to the phonon excitation spectrum of the spherical shape except that the $0_{3}^{+}$ state is repelled to a higher energy level [18, 19]. Thus at the previous three-phonon level, the $0_{3}^{+}$ state does not exist at all [20]. The B(E2) values among the normal states can be well explained in this model. In particular, it was recently found that the anomalous evolution trend of the electric quadrupole moment of the $2_{1}^{+}$ state of 108-116Cd can be explained by this model, which implies that the SU3-IBM is the valuable attempt to resolve the Cd puzzle [20].
Here, there exist two important conclusions: (1) higher-order interactions are important and the third-order interaction can be even dominant [18, 24, 26], (2) the SU(3) symmetry governs all the interactions except for the pairing force (the $d$ boson number operator). These imply that the idea of the mean field requires some corrections, meaning that the two-body interaction may not necessarily dominate the residual ones. And an important question is, where does this SU(3) symmetry in the SU3-IBM come from?
In previous IBM, the SU(3) symmetry is only one reduction limit of the SU(6) group [21], so it is an accidental product of the shape symmetry spontaneous breaking of the nucleus (from the spherical shape to the prolate shape) [27]. Its appearance does not require an explanation. But in the SU3-IBM, this situation becomes different, because all the quadrupole shapes are described by the SU(3) symmetry, so a deeper understanding is needed.
Obviously, the success of the SU3-IBM proves that the key to the Cd puzzle, the fact that the spherical nucleus does not exist, is related to the SU(3) symmetry. This result is very interesting. In the SU(3) representation, (0,0) is the spherical shape. Thus the Cd puzzle implies that the shape of the nucleus can not be the (0,0) representation of the SU(3) group. In this Letter, a simple explanation on this impossibility is given.
The Cd puzzle is the low energy phenomenon of the nucleus. It is hard to imagine that this phenomenon may be related to the high energy behavior. Here some unexpected relations are found.
A Nucleus is made of protons and neutrons, which are collectively called nucleons. The nucleons are also composite particles, consisting of valence quarks, sea quarks, and gluons. Considering that the binding energy of the nucleus is less than one percent of the nuclear energy, many researchers would think that the nucleons in the nucleus are the same as the free nucleons, or treat them as point-like particles. However, the EMC effect found in 1982 rejected this perspective [28, 29, 30, 31].
The EMC effect illustrates that the result of deep inelastic scattering is related to the target of the used nucleus, and the effect increases as the mass number of the nucleus increases [28, 29, 30, 31]. This means that the quark-gluon distribution of the nucleons in the nucleus is not the same as the ones in the free nucleons. Recently, researchers have generally agreed that this phenomenon is related to the medium modification of the nucleus, fundamentally, to the color confinement [31]. This means that in the nucleus, except for the nucleons and various instantaneous mesons, quarks and gluons need to be considered. In the nucleus, quarks and gluons are not only bound within the nucleons, but also occasionally appear between the nucleons. Thus the nucleons in the nucleus may not be temporarily color-neutral.
Researchers have also studied the effect of the quarks and gluons in the nucleus [32, 33, 34, 35]. Just as Cloët et al. said:“ it is unlikely that the nucleon-meson based approaches can remain valid or contain the correct degrees of freedom for all processes at all energy scales.” The quark-gluon degree of freedom is needed in the discussion of nuclei. The key is to find clear or revelatory evidence. It’s hard to imagine that quarks and gluons are only confined to the interior of the nucleons and mesons while the nucleus has a larger volume.
Quarks and gluons have color degrees of freedom which has SU(3) symmetry. The color confinement is a conjecture that has not been proved by theory. It means that composite particles made up of quarks and gluons, such as hadrons or nuclei, are color-neutral, or the (0,0) representations of SU(3) group. Here I focus on the nuclei. Hydrogen nucleus is the proton, so it can be expressed by the (0,0) representation of the SU(3) group. The same is true for the neutron, and also true when extended to nuclei with larger mass number $A$. The whole composition of the quarks and gluons must be color-neutral.
The EMC effect means that the nucleons in the nucleus are different from the free nucleons, which are the true embodiment of the color confinement. Although the specific physical mechanism is unknown, if the color confinement in the free nucleons is broken, it is clear that these nucleons in the nucleus cannot be represented by the (0,0) representation of the SU(3) group. As long as the color degree of freedom appears between the nucleons in the nucleus, this conclusion is inevitable. It should also be emphasized that this color effect in nuclear medium becomes more pronounced as the number of nucleons increases.
This is a bold step forward understanding the color degree of freedom in the nucleus in this Letter. Although this is to be expected, there is a lack of evidence to suggest this in previous studies.
Thus a hypothesis can be given for a nucleus, that only the nucleus itself is a trivial (0,0) representation of the SU(3) group. This means that the nucleons in the nucleus, and their combinations, as long as they are not the nucleus, cannot be the (0,0) representation of the SU(3) symmetry, no matter how complex the specific interactions between them are. Because if some of the nucleons are combined into the (0,0) representation, implying that this is a nucleus, it means that the original nucleus has been fission.
This assumption certainly needs the theoretical confirmation (color confinement), but here we will see that it directly gives the conclusion that a spherical nucleus does not exist. It has now been confirmed that the nuclear force can be divided into three parts: the mean field, short-range strong correlation, and long-range correlation [31]. Recent experiments have also demonstrated that the EMC effects are associated with short-range strong correlations, but the more comprehensive correlations with the three parts remains unclear. It is already known that the EMC effects can be related to the corrections of the mean-field nucleons and short-range strongly correlated nucleons, but the exact mechanism is still unclear [31, 36, 37].
Thus some researchers will doubt whether such medium correction will be related with the low-energy nuclear structure. Recently Machleidt and Sammarruca said:“ it is a good question to ask whether medium modifications of nuclear forces show up in a noticeable way and/or are even needed for quantitative nuclear structure predictions.” and asked: “Are the energies typically involved in conventional nuclear structure physics low enough to treat nucleons as structure-less objects?” Here I give the answer, even at the low-lying excitation, it is not possible. The Cd puzzle means the nucleon modifications and the color degree freedom is also needed if the assumption is correct.
This explanation is very simple. If the representation of the nucleus itself is the (0,0) representation of the SU(3) group, then no part of it can be the (0,0) representation, which means that the mean field cannot be the (0,0) representation, and the long-range correlation is not the (0,0) representation too. Although this conclusion is very simple, it is strict.
This requires a further detailed explanation. Since the representation of the final nucleus is (0,0), the product of the representation of all nucleons is (0,0), which has many possibilities. All possibilities need to be considered to understand the nature of the nucleus. Each possibility can be expressed as
$(\lambda_{1},\mu_{1})\otimes(\lambda_{2},\mu_{2})\otimes\cdot\cdot\cdot\otimes(\lambda_{A},\mu_{A})=(0,0),$ (1)
here $A$ is the nucleon number, and $(\lambda_{k},\mu_{k})$ is the SU(3) representation of the kth nucleon, $1\leq k\leq A$. This is actually the expression of the color confinement in the nucleus. What is the $(\lambda_{k},\mu_{k})$ is still unknown, which needs further determined. But it can not be (0,0). In the simple case they can be considered as (1,0) and (0,1). For example, the deuteron can be expressed as
$(1,0)\otimes(0,1)=(0,0),$ (2)
and the short-range correlation pair can be expressed as
$(1,0)\otimes(0,1)=(1,1).$ (3)
If we want to know the proportion of the short-range strong correlations in the nucleus with $A$ nucleons, the number of equation (3) in all possibilities can be counted, which will be discussed in the next paper. It should be stressed here that the intermediate coupling result can not appear as the (0,0) representation. When $A$ increases many coupling possibilities can be obtained. All these possibilities will be divided into the mean field, short-range correlations and long-range correlations. Therefore, these three parts are related to each other, and they cannot be expressed by (0,0). Although the way of the division is still unclear, the result is clear.
It should be emphasized that not one possibility corresponds to one part, but the three parts are possible. It is not clear how to translate into the three parts, perhaps requiring support from experimental data. The influence of the color confinement in the nucleus is emphasized here, especially establishing the correlation between the low and high energy behaviors. Every possibility of the nuclear color confinement is to be considered, so all possibilities must affect the properties of the nucleus as a whole.
It is also important to emphasize that, although the color SU(3) symmetry is used in the discussion, the specific color state is not discussed here. The (0,0) represents a color-neutral state, but other $(\lambda_{k},\mu_{k})$ are just a phenomenological representation. This does mean that quarks or gluons are temporarily transferred from one nucleon to another, but the exact mechanism is not clear. Here, it is only emphasized that quarks and gluons cannot be fully bound within the nucleons in the nucleus. When approximately dividing into the mean field, short-rangle strong correlations and long-range correlations, although we still use the SU(3) symmetry, this may be an approximation and may have new physical significance, such as shape, because this is a collective effect. If the mean field is $(\lambda_{m},\mu_{m})$, the short-range strong correlation is $(\lambda_{s},\mu_{s})$, the long-range correlation is $(\lambda_{l},\mu_{l})$, the division can be expressed as
$(\lambda_{m},\mu_{m})\otimes(\lambda_{s},\mu_{s})\otimes(\lambda_{l},\mu_{l})=(0,0).$ (4)
Finding the division law is at the heart of the further research.
First the mean filed is not (0,0). Although the magic number can be explained by the spherical single-particle potential, a potential field that does not deviate too much from the sphere is also possible. In the experiments, some researchers have also begun to find that some magic number nuclei exist a collective excitation [39, 40, 41, 42], so they are not spherical. Just as Garrett et al. said:“ Rather than proceeding from spherical, closed-shell nuclei through a region of spherical vibrators before encountering deformation, deformation and shape coexistence may be confronted immediately, even at the closed shell.” [15] So the conclusion is credible. This result can also help us to find better single-particle potentials to understand the low-energy properties of nuclei.
Second the long-range correlations are not (0,0). This long-range correlations lead to the deformations of the nucleus. If they are not (0,0), it means that it cannot be spherical. This is exactly what the Cd puzzle is trying to reveal. This conclusion is also supported when the Cd puzzle is verified by the SU3-IBM. In the new spherical-like $\gamma$-soft mode, the SU(3) third-order Casimir operator is as important as the second-order one. This is a new kind of deformation. This means that when the valence nucleons increase, the two-body residual interaction does not dominate. This is not quite the usual mean field viewpoint. I guess that the multiple possibilities of the color confinement in the nucleus in equation (1) let the low-energy two-body interactions interfere destructively, reducing it.
These are the main conclusions to be given in this Letter.
The deformation of the nucleus is a fundamental problem of nuclear structure. In the previous understanding, because the magic number nucleus is spherical, the deformation needs to occur when the number of valence nucleons is relatively large. The Cd puzzle [4, 17, 5] and the energy spectra of the magic nuclei [39, 40, 41, 42] do not support such a conclusion. If the mean field is not spherical, then new collective excitation can emerge when the valence nucleons appear. If we can get the explicit expressions of $(\lambda_{m},\mu_{m})$ and $(\lambda_{l},\mu_{l})$, we will understand exactly what happens.
This hypothesis is strongly supported by the EMC effect, which needs to be further verified. The Cd puzzle is the direct consequence of this hypothesis. Although this explanation is easy, it fits with the results of the experiments. The conclusion that spherical nucleus does not exist can be confirmed by the Cd puzzle experiments and can be supported by other surrounding nuclei, such as Te, Pd nuclei [5]. The results of these experiments can be well described by the SU3-IBM [18, 19, 26], which further supports this hypothesis.
The SU3-IBM explains not only the normal states of the Cd nuclei, but also the B(E2) anomaly phenomenon [43, 44, 45], which cannot be explained by previous theories of nuclear structure [47, 46, 48, 49]. This model is also used to describe the prolate-oblate asymmetric shape evolution, which reveals the deformation of the prolate shape is nearly twice of the one of the oblate shape [24]. This model can explain the properties of 196Pt at a better level [20, 50] and also the E(5)-like behavior of 82Kr [25]. All these mean that the SU3-IBM is able to give a more accurate description of the low energy excitations of the nuclei and is a more efficient model. So the conclusion that the spherical nucleus does not exist is plausible, which supports the correctness of the hypothesis.
The result is somewhat unexpected. For the first time, it explicitly points out a clear correlation between the low and high energy behaviors of the nucleus, and while the true reason for the correlation is still unknown, it does exist. Of course, it can be guessed that this correlation comes from the quark-gluon effect in the nucleus, which would otherwise not explain the SU(3) symmetry in the SU3-IBM. These results illustrate that the color degree of freedom is influential even in the field of low-energy nuclear structure. This can be supported by the SU(3) symmetry of the SU3-IBM.
So this hypothesis is supported by both experiment and theory. The role of the quarks and gluons in the nucleus, though not clear, exists and works at low energy excitation. This has been overlooked in previous studies. Of course, this is inevitable. Because the change of the quarks and gluons in the nucleus is the behaviors found in the high-energy region, it is difficult to expect that the EMC effect can affect the low-energy nuclear structure. This requires support from the experiments. The Cd puzzle, as well as its SU3-IBM description, becomes the possible evidence.
Since the EMC effect increases with the number of nucleons, it can be guessed that the effect of this color degree of freedom is more obvious for heavy nuclei. The fact is true. The SU3-IBM explains the prolate-oblate asymmetric phase transition in the Hf-Hg region, where the SU(3) third-order Casimir operator describes the oblate shape. The recent discovery of the boson number odd-even effect in 196-204Hg proves the correctness of this theory [26], and also indicates the emergence of previously completely unexpected phenomena in heavy nuclei. These strange color-related phenomena may be very difficult to detect in light nuclei. It is not appropriate to draw conclusions from mere experiences with light nuclei. Thus the influence of the color confinement in the nucleus should be found in the nucleus with more nucleons, especially the medium-heavy nuclei.
How to understand the color confinement in the nucleus is important for the following studies. Equation (1) is the starting point of the discussions. In the simplest case, the representations of the nucleons are (1,0) and (0,1). For more complex situations, (2,0) (0,2) and (1,1) will also need to be considered. The key thing is how to extract the mean field, short-range strong correlations and long-range correlations from all possibilities of the color confinement, the equation (4). This is an interesting and urgent question. It should be noticed that if this division rule exists, it must be sensitive to the nucleon number $A$. We expect that this may have something to do with the magic numbers and the reduction of the SU(3) second-order interaction.
Currently, the main tool for nuclear force researches is the effective field theory [38, 51, 52, 53]. In their discussions, the color degree of freedom is not considered, which means that the quarks and gluons are simply bound to the nucleons. The color confinement in the nucleus is not considered in these studies. It can be seen from this Letter that for a complete understanding of the nuclear force, even at low energy, quarks and gluons need to be considered. How to combine the idea of the effective field with equation (1) is an important work. Just as Miller said:“ all three scales (the nuclear size, the internucleon separation distance and the nucleon size) are must be understood to truly understand the physics of nuclei.” [35] If the division equation (4) exists, the idea of the effective field theory is still applicable, only adding a new limitation.
In conclusion, a simple explanation for the absence of the spherical nuclei is given based on the EMC effect, short-range strong correlations, the Cd puzzle and other new experimental results, and the calculation of the SU3-IBM. It can be obtained that the Cd puzzle is the result of the color confinement in the nucleus. It is clearly pointed out that the color degree of freedom is necessary even in low energy excitation of nuclear structure. The correlations between the low energy and high energy behaviors in the nucleus does exist, and this is the first time to be found here, but what it is requires more further studies.
Here a new theoretical framework to understand nuclear forces is established phenomenologically. Since this explanation is very simple and rigorous, the new connection deserves serious consideration. It does not conflict with previous effective field theories. The possibility of the color degree of freedom between the nucleons is considered here. The SU3-IBM supports this possibility and gives a guide for future theoretical developments. In future, the possible results of this new perspective will be discussed.